OpenStudy (anonymous):

Please Help ASAP!! Pleaseee No one has helped me with this one!! Which equation best represents the data in the table below? x 0 2 4 6 8 y 7 3 -1 -5 -9 2x + y = 8 x + 2y = 7 x + 2y = 8 2x + y = 7

OpenStudy (texaschic101):

pick a set of points on your table....(0,7) x = 0 and y = 7 now sub those points into your answer choices and see which one it satisfies 2x + y = 8 2(0) + 7 = 8 7 = 8 (nope, not this one) x + 2y = 7 0 + 2(7) = 7 14 = 7 (nope, not this one either) x + 2y = 8 0 + 2(7) = 8 14 = 8 (nope) 2x + y = 7 2(0) + 7 = 7 7 = 7 (BINGO...found it) you can even check it with other points from your table.. (2,3) x = 2 and y = 3 2x + y = 7 2(2) + 3 = 7 4 + 3 = 7 7 = 7 (correct) you will find, that whatever points you pick from your table, will satisfy the equation 2x + y = 7.....which is your answer